OpenStudy (anonymous):

calculate the simple interest and maturity value for the following problem principal $18,000 interest 41/2 time is 18 months

OpenStudy (anonymous):

I = P r t = 18,000* 4.5% * 18 = $14,580 V= P (1+r)^18 = 18,000 ( 1.045)^18 = $39,753

OpenStudy (anonymous):

I thought 4 1/2% was .045

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

where do you don't understand?

OpenStudy (anonymous):

you did 4 1/2 as 4.5 and theanswer seems to much for the interest

OpenStudy (anonymous):

Why don't you calculate yourself?

OpenStudy (anonymous):

dont be nasty god is good in my life all I said was i thought 4 1/2 % was .045

OpenStudy (anonymous):

That's why I said you need to do it yourself to confirm it's .045!

OpenStudy (anonymous):

ok thank you i see you did the problem but whats the maturity of the loan